OpenStudy (anonymous):

whats the equation for the sum of the series 1/(n^2+i^2)?

OpenStudy (anonymous):

Is that a double sum, like \[\sum_{n=1}^{\infty}\sum_{i=1}^{\infty} \frac{1}{n^2 + i^2}\] or is that something else?